Find the two consecutive integers whose product exceeds their sum by 19.
How can I solve this? What would the exquation be??
I'm really having a hard time finding the equations. Please help.

Answer by edjones(8007)

Let integers be x, x+1
x(x+1)=x+x+1+19
x^2+x=2x+20
x^2-x-20=0 subtract 20 and 2x fr4om each side.
(x-5)(x+4)=0 Factor.
x=5, x=-4
Check:
5*6=5+6+19...30=30 true
-4*-3=-4-3+19...12=12 true
So there are 2 sets of correct answers 5,6 and -4, -3.
Ed